OpenStudy (anonymous):

Block m1 of mass 2m and velocity v0 is trav- eling to the right (+x) and makes an elastic head-on collision with block m2 of mass m and velocity −2 v0 (i.e., traveling to the left). What is the velocity v1 0 of block m1 after the collision?

OpenStudy (anonymous):

2m(vo)+m(-vo)=v(m+m)

OpenStudy (anonymous):

would it be v1=+v0

OpenStudy (anonymous):

The important part here is that although their momenta are equal and opposite, the kinetic energies are not. You need to set up equations for the momentum conservation and the kinetic energy conservation. A tentative answer (i've not been too careful with my signs in the algebra... :/_ is v1=-0.369v0 (i.e. it has bounced backawrd from original direction). Good luck!

OpenStudy (anonymous):

not an option for the answers im provided with

OpenStudy (anonymous):

what is the answer you have? and did the actual physics I explained make sense?

OpenStudy (anonymous):

1. v10 = −3/2v0 2. v10 = −v0 3. v10 =3/2vo 4. v10 = 3 v0 5. v10 = 2 v0 6. v10 = v0 7. v10 = −3 v0 8. v10 = −2 v0

OpenStudy (anonymous):

I know you have to use the momentum conservation equation and the velocity equation for ellastic collisions

OpenStudy (anonymous):

take a look at this http://hyperphysics.phy-astr.gsu.edu/hbase/elacol2.html that is basically the physics I used but I may well have done something wrong along the way..

OpenStudy (anonymous):

that way i get m/3m(vo)

OpenStudy (anonymous):

that be 1/3 vo and its not an option :(

OpenStudy (anonymous):

oh i see the problem in them examples on the website the other object isnt moving in the opposite direction. it my problem its like two cars driving straight toward another and crashing. that makes the second velocity negative

OpenStudy (anonymous):

i think you take v1 initial - v2 initial= v2 final - v1 final solve for one of the final velocitys then plug that back into the conservation of momentum to solve for the other final velocity

OpenStudy (anonymous):

I had another go at this, I now get v10 = -v0 from solving conservation of momentum and kinetic energy equations, i.e the block bounces back: with the same speed in the opposite direction. Using the website link equation, the second body is needed to start at rest. this has effect that the initial velocity is 3 times larger for block 1 (i.e. vo is now 3vo). Put that in and you get v10 = v0, but no minus sign..